OpenStudy (anonymous):

what am i'm doing wrong with this problem to get this answer? 0.96y − 0.79 ≤ 0.21y + 0.46 = 0.96y - 0.21y – 0.79 < 0. 21y - 0.21y + 0.46 =0.04y < 0.46 = 0.04y/0.04 < 0.46/0.04 = y < 0.46/0.04 = y < 11.5 Check: 0.96y − 0.79 ≤ 0.21y + 0.46 Using 0 0.96() – 0.79 < 0.21(0) + 0.46 0.79 < 0.46

OpenStudy (anonymous):

0.96y − 0.79 ≤ 0.21y + 0.46 = 0.96y - 0.21y – 0.79 < 0. 21y - 0.21y + 0.46 =0.04y < 0.46 = 0.04y/0.04 < 0.46/0.04 = y < 0.46/0.04 = y < 11.5 Check: 0.96y − 0.79 ≤ 0.21y + 0.46 Using 0 0.96(0) – 0.79 < 0.21(0) + 0.46 0.79 < 0.46

OpenStudy (anonymous):

=0.04y < 0.46 (How did you get this?)

OpenStudy (anonymous):

Third line should read: .75y - 0.79 < 0.46

OpenStudy (anonymous):

Most students don't like working with decimals. So, it's best to get rid of them by multiplying the entire equation by an appropriate factor of 10, in this case, 100.

OpenStudy (anonymous):

96y - 79 <= 21y + 46 75y <= 125 y <= 125/75 y <= 5/3

OpenStudy (anonymous):

If you need an answer in the form of a decimal then 5/3 = 1.66666...

OpenStudy (anonymous):

It looks like you went like this: \[.75y -.79 < .46 \Rightarrow .04y < .46\] but that is incorrect. You cant subtract the .75y and .79, they are not like terms. This is what you should have done: \[.75y-.79+.79 < .46 + .79 \Rightarrow .75y < 1.25\] then continue on from there :)

OpenStudy (anonymous):

thanks u both for your help I kind of understand it a lil better.